A and X are the matrices
a b
c d
and
x y
u v
respectively, where b is not equal to zero. Prove that if AX = XA then u = cy/b and v = x + (d − a)y/b. Hence prove
that if AX = XA then there are numbers p and q such that X = pA + qI, and find
p and q in terms of a, b, x, y.
1
Expert's answer
2020-06-01T19:07:48-0400

A="\begin{bmatrix}
 a & b \\
 c & d
\end{bmatrix}" and X="\begin{bmatrix}
 x & y \\
 u & v
\end{bmatrix}" and b is not equal to zero.

now given,

AX = XA

or "\begin{bmatrix}
 a & b \\
 c & d
\end{bmatrix}\begin{bmatrix}
 x & y \\
 u & v
\end{bmatrix}" = "\begin{bmatrix}
 x & y \\
 u & v
\end{bmatrix}\begin{bmatrix}
 a & b \\
 c & d
\end{bmatrix}"


or "\begin{bmatrix}
 ax+bu & ay+bv \\
 cx+du & cy+dv
\end{bmatrix}" = "\begin{bmatrix}
 ax+cy & bx+dy \\
 au+cv & bu+dv
\end{bmatrix}"


As the two matrices are equal we can compare the corresponding elements of the two matrices.


"\therefore" "ax+bu = ax+cy"

or "bu=cy"

or "u = cy\/b" ( as b is not equal to zero)

(proved)

Again,


or "ay+bv = bx+dy"

or "bv = dy-ay+bx"

or "bv = (d-a)y+bx"

or "v= (d-a)y\/b+x" (proved)


Now

or "X=pA+qI"

or "\begin{bmatrix}
 x & y \\
 u & v
\end{bmatrix}" = "p\begin{bmatrix}
 a & b \\
 c & d
\end{bmatrix}" + "q" "\begin{bmatrix}
 1 & 0\\
 0 & 1
\end{bmatrix}"

or "\begin{bmatrix}
 x & y \\
 u & v
\end{bmatrix}" = "\begin{bmatrix}
 p a &pb \\
 pc & pd
\end{bmatrix}" + "\begin{bmatrix}
 q & 0 \\
 0 & q
\end{bmatrix}"

or "\begin{bmatrix}
 x & y \\
 u & v
\end{bmatrix}" = "\begin{bmatrix}
 pa+q &p b \\
 pc & pd+q
\end{bmatrix}"


As the two matrices are equal we can compare the corresponding elements of two matrices.


"\therefore" "x=pa+q" and "y= pb"

Now "y=pb"

or "p = y\/b" . Answer


again

"x=pa+q"

or "x= (ay)\/b+q"

or "q= x-((ay)\/b)" Answer